OpenStudy (amazingkenma):

6x-10=6x+10 Find the value of x A) 15 B) 26 C) 46 D) 80

OpenStudy (eliesaab):

Are you sure you typed your question right?

OpenStudy (eliesaab):

As it is, it reduces to 0 x=20 which does not have any solution

OpenStudy (maxwellfish):

16x-10=6x+10 X=2

OpenStudy (sshayer):

as it is typed 0=20 which is impossible. 6x-10=6x+10 6x-6x=10+10 0=20